# On stabilizers in finite permutation groups

### v2

@@ 1. Introduction
In this paper, we are interested in a small variation on these questions. More precisely, we fix only the bare minimum (so one subset, or a pair of points) and we do not aim to find a trivial stabilizer, since this would be impossible in general, but to find a stabilizer that has at least some good properties. We will see that this can be done in a few interesting situations. Our first main result is the following:
@@
Let $G\leqslant\mathop{\mathrm{Sym}}(\Omega)$ be a solvable permutation group. Then there exists a subset of $\Omega$ whose setwise stabilizer has all orbits of length at most $6$.
@@
For primitive solvable groups, we can find a subset whose stabilizer has orbits of length at most $3$ and is metabelian. In the beginning we hoped to prove the same for all transitive solvable groups, but this happened to be false (see Examples 2.8 and 2.9). In fact, the results above are the best possible, and answer two questions of Babai [Bab22, Questions 7a-7b]. Theorem 1.1 seems to be the first statement concerning stabilizers of imprimitive groups where the property involved is not preserved by group extensions, and part of the difficulty of the proof is in finding the right statement to prove by induction. [^1] Babai [Bab22, Theorem 12.1] proves the same result with a nonexplicit constant.
We come to the question of fixing two points. Extending the work of Seress [Ser96], Burness [Bur21] showed that in a primitive group with solvable stabilizer, there always exist $5$ points whose pointwise stabilizer is trivial. For primitive solvable groups, in [LS24] the authors observed that there exist two points whose pointwise stabilizer has bounded derived length. On the other hand, motivated by various applications, Li and Zhang [LZ11] have classified the solvable maximal subgroups of the almost simple groups (see also [Bur21]). Combining their work with some properties of the finite simple groups, we remark the following:
@@
Let $G$ be an almost simple group, and let $H<G$ be a maximal subgroup. If $H$ is solvable, then the derived length of $H$ is at most $10$. The constant $10$ is realized only by the solvable maximal subgroups of the Fischer group $Fi_{{}_{23}}$.
@@
The proof of Theorem 1.4 is an application of Corollary 1.2, and it is perhaps surprising that, to answer a question about pointwise stabilizers of primitive nonsolvable groups, one has to deal with setwise stabilizers of arbitrary solvable permutation groups. The idea, usually applied in the context of imprimitive linear groups, is using a partition to construct a vector with small centralizer (see [MW04, Sec. 4]).
The paper is organized as follows. In Section 2 we study setswise stabilizers, proving Theorem 1.1 and other related results (we also answer [Bab22, Question 7c], and give an elementary proof of a theorem of Cameron, Neumann and Saxl). In Section 3, we prove Theorems 1.3 and 1.4. The methods are a mixture of group theory and combinatorics.
Acknowledgments: I am grateful to S. Eberhard for many useful conversations. After the public announcement of a first version of this paper, I received precious comments from D. Gluck, H.Y. Huang and S.V. Skresanov. The first asked for a result in the spirit of Corollary 2.11. The second noticed that the constant $10$ in Theorem 1.3 is attained by $Fi_{{}_{23}}$. The third pointed out Babai’s paper [Bab22], and helped me with the computer calculations in Lemma 2.6. Finally, I am thankful to L. Pyber for the many ideas he has given me in these years.
2. Setwise stabilizers of solvable groups
@@
When $G$ acts on a finite set $\Omega$, then $G$ acts naturally on the power set $\mathcal{P}(\Omega)$. If $\Delta\subseteq\Omega$, we write $C_{G}(\Delta)$ for the pointwise stabilizer, and $Stab_{G}(\Delta)$ for the setwise stabilizer. As usual, we say that $\Delta$ lies in a regular orbit if $Stab_{G}(\Delta)=1$. We say that $\Delta$ lies in a binary orbit if all the orbits of $Stab_{G}(\Delta)$ on $\Omega$ have length at most $2$, in a ternary orbit if all the orbits of $Stab_{G}(\Delta)$ on $\Omega$ have length at most $3$, and in a good orbit if all the orbits of $Stab_{G}(\Delta)$ on $\Omega$ have length at most $6$. Of course,
@@
We point out that $f(b)=O(\log b)$ [Dix68]. When combined with [Ser96], this gives that in a primitive solvable group there exists a setwise stabilizer of derived length at most $f(4)=3$. We are going to see that a bit more is true.
@@ 2.1. Primitive solvable groups
Similarly, we remark the following:
Lemma 2.3.
Proof.
As before, using [MW04, Lemma 4.1(ii)]. ∎
$\dfrac{1^{1}}{1_{1_{1}}}$ GAP Id Regular Orbits on $\mathcal{P}(\Omega)$ Binary Orbits on $\mathcal{P}(\Omega)$ Ternary Orbits on $\mathcal{P}(\Omega)$ Good Orbits on $\mathcal{P}(\Omega)$
$\ell_{1,2}$ $\ell_{2,2}$ $\ell_{3,2}$ $\ell_{6,2}$
@@
As an example to help reading Table 1, we remark that PrimitiveGroup(4,2) is $\mathop{\mathrm{Sym}}(4)$. There is no regular orbit of $\mathop{\mathrm{Sym}}(4)$ on $\mathcal{P}(\{1,2,3,4\})$, the stabilizer of $\{1,2\}$ has two orbits of length $2$, and the stabilizers of $\{1\}$ and $\{1,2,3\}$ have orbits of length $3$. We also remark that PrimitiveGroup(8,2) is $\mathrm{A\Gamma L}_{1}(8)$, while PrimitiveGroup(9,7) is $\mathop{\mathrm{AGL}}_{2}(3)$.
We now report the following sharp result:
@@
See [MW04, Lemma 4.1(iii)]. ∎
2.2. Imprimitive groups, colorings
The standard argument to deal with imprimitive groups is [Glu83, Theorem 2], which was used by Gluck to show that permutation groups of odd order have a regular orbit on the power set. This does not work with a property such as “have all orbits of length at most $C$”, where $C$ is a constant. We use a different method based on good orbits and partitions that we are going to explain.
The distinguishing number of a permutation group $G$ on $\Omega$ is the smallest number of parts in a partition of $\Omega$ with the property that only the identity fixes every part. For example, this is $2$ if and only if $G$ has a regular orbit on the power set. For more details, see [BC11, Sec. 2.2]. We refine the notion as follows:
@@ 2.2. Imprimitive groups, colorings
Let $G\leqslant\mathop{\mathrm{Sym}}(\Omega)$, and $k\geq 2$. A $k$-coloring is a coloring of $\Omega$ with at most $k$ colors $\{1,\ldots,k\}$. A regular $k$-coloring is a $k$-coloring such that the identity is the only element of $G$ that fixes every color. A binary $k$-coloring is a $k$-coloring such that the stabilizer of the coloring in $G$ has all orbits of length at most $2$. Ternary $k$-colorings and good $k$-colorings are defined accordingly.
The total number of $k$-colorings of $\Omega$ is $k^{|\Omega|}$, but many of these are equivalent under the action of $G$. Fixed $G\leqslant\mathop{\mathrm{Sym}}(\Omega)$, for each $k\geq 2$ we write
@@
for the numbers of $G$-orbits of regular, binary, ternary, and good $k$-colorings. It is obvious that, for every $i\leq j$ and $k$,
(2.1) $\ell_{i,k+1}\geq\ell_{i,k},\hskip 28.45274pt\mbox{ and }\hskip 28.45274pt\ell_{j,k}\geq\ell_{i,k}.$
We observe that describing a $2$-coloring is equivalent to describing a subset. Therefore, $\ell_{1,2}$, $\ell_{2,2}$, $\ell_{3,2}$ and $\ell_{6,2}$ coincide respectively with the numbers of regular, binary, ternary and good orbits on the power set. Restated in terms of colorings, Lemma 2.4 says that $\ell_{1,4}\geq 1$ for every primitive solvable group (one color is needed for the complement).
A portion of the proof of Theorem 1.1 relies on showing that $\ell_{i,6-i}\geq 5$ for all primitive solvable groups and each $i=1,2,3$. The following combinatorial lemma is useful to estimate $\ell_{i,k}$ for large $k$.
@@
Fix $i\in\{1,2,3,6\}$. If $\ell_{i,2}\geq 1$, then $\ell_{i,k}\geq{k\choose 2}$. If $\ell_{i,j}>\ell_{i,j-1}$ for some $3\leq j\leq k$, then $\ell_{i,k}\geq{k\choose j}$.
@@
The inequality $\ell_{i,j}>\ell_{i,j-1}$ means that there exists a suitable coloring that involves precisely $j$ colors. Now suppose we have such a $j$-coloring. For each subset of cardinality $j$ of the $k$ available colors, we obtain an inequivalent suitable $j$-coloring. ∎
The following result requires a slightly deeper investigation. In the notation $\ell_{i,k}$, we write $\ell_{i,k}(G)$ to specify the group we are referring to.
@@
Let $A$ be a primitive solvable group. Then $\ell_{2,3}(A)\geq 5$.
@@
Let $\Omega$ be the base set. By Lemma 2.2 and (2.1), we can assume $|\Omega|\leq 9$. Since we are interested in a lower bound on the number of orbits, it is enough to deal with the groups
@@
It is easy to see that $\ell_{2,3}$ is at least $5$ in all these cases. In particular, using [GAP4], the command
@@
We prove Theorem 1.1 through the following stronger statement, that lends itself to an inductive argument.
@@
We work by induction on the cardinality of $\Omega$. Suppose that $\Omega_{1},\ldots,\Omega_{t}$ are the orbits of $G$ on $\Omega,$ and that we can find subsets $\Delta_{j}\subseteq\Omega_{j}$ such that $Stab_{G}(\Delta_{j})$ has orbits of length at most $6$ on $\Omega_{j}$, for each $j=1,\ldots,t$. Then $Stab_{G}(\sqcup_{j}\>\Delta_{j})$ will surely work. It follows that we can suppose that $G$ is transitive. Primitive groups can be handled with Lemma 2.2 and Table 1. Let $\Omega=\Omega_{1}\times\Omega_{2}$, with $|\Omega_{1}|=a\geq 2$. Let $G\leqslant B\wr_{a}A$, with $A$ primitive on $\Omega_{1}$, and $B$ transitive on $\Omega_{2}$. By the inductive hypothesis applied to $\Omega_{2}$, we are in one of the following cases:
@@
    We are going to use the good orbits of $B$ on $\mathcal{P}(\Omega_{2})$, say $\Sigma_{1}^{B},\ldots,\Sigma_{5}^{B}$, as colors for $\Omega_{1}$. Indeed, for each regular $5$-coloring of $\Omega_{1}$, say $\Omega_{1}=\sqcup_{j=1}^{5}\Delta_{j}$, we can construct a good orbit $\Delta^{G}$ for $G$ as follows:
    $\Delta=\{(x,y)\in\Omega_{1}\times\Omega_{2}:x\in\Delta_{j}\mbox{ and }y\in\Sigma_{j}\mbox{ for some fixed }j\}.$
    The setwise stabilizer of $\Delta$ in $G$ can be embedded in the semidirect product between the direct product of the stabilizers of the $\Sigma_{j}$’s in $B$ (which have orbits of length at most $6$), and the stabilizer of the regular $5$-coloring in $A$ (which is trivial). Inequivalent $5$-colorings of $\Omega_{1}$ provide inequivalent good orbits of $\Omega$. In particular, there are at least $\ell_{1,5}(A)$ good orbits of $G$. Now $\ell_{1,4}(A)\geq 1$ by Lemma 2.4, and so, from Lemma 2.5,
    $\ell_{1,5}(A)\geq\min_{j=2,3,4}{5\choose j}=5.$
@@
    Here $\ell_{2,2}(B)=3$. Using the binary orbits of $B$, and the ternary $3$-colorings of $\Omega_{1}$, as before we can construct $\ell_{3,3}(A)$ good orbits of $G$. In fact, the length of any orbit will be bounded by $2\cdot 3=6$. Now $\ell_{3,3}(A)\geq\ell_{2,3}(A)\geq 5$ by (2.1) and Lemma 2.6.
@@
    We can assume $B\cong\mathop{\mathrm{Sym}}(3)$, so that $\ell_{3,2}(B)=4$. Using the ternary orbits of $B$, and the binary $4$-colorings of $\Omega_{1}$, as before we can construct $\ell_{2,4}(A)$ good orbits of $G$. In fact, the length of any orbit will be bounded by $3\cdot 2=6$. Now $\ell_{2,4}(A)\geq\ell_{2,3}(A)\geq 5$ by (2.1) and Lemma 2.6. ∎
The cardinality of $\Delta$ in Theorem 1.1 grows linearly with the cardinality of $\Omega$, which is necessary as shown by $B\wr K$ when $B$ has some orbit of length exceeding $6$. The bound on the derived length follows easily:
@@
The next example provides a transitive solvable group where no subset gives a metabelian stabilizer.
@@
Let $B=\mathop{\mathrm{Sym}}(4)\wr_{4}\mathop{\mathrm{Sym}}(4)$ act imprimitively on $\Omega=\{a,b,c,d\}\times\{1,2,3,4\}$. It is not hard to see that $B$ has a unique orbit on the power set whose setwise stabilizer is metabelian, namely $\Delta^{B}$, where
@@
The setwise stabilizer $H$ of $\Delta$ in $B$ is isomorphic to $\mathop{\mathrm{Sym}}(3)\times\mathop{\mathrm{Sym}}(3)\times D_{8}$. Now let $G=B\wr_{2}\mathop{\mathrm{Sym}}(2)$ act on $\Omega\times\{\alpha,\beta\}$. If we color both $\Omega\times\{\alpha\}$ and $\Omega\times\{\beta\}$ using $\Delta$, then the resulting setwise stabilizer will be isomorphic to $H\wr_{2}\mathop{\mathrm{Sym}}(2)$, whose derived length is $3$. Otherwise, we are forced to color a subset, say of $\Omega\times\{\alpha\}$, which does not provide a metabelian setwise stabilizer in $B$. So there is no hope of getting a metabelian setwise stabilizer in $G$.
The next similar example involves $\mathrm{A\Gamma L}_{1}(8)$, and provides a transitive solvable group where no subset gives a stabilizer with all orbits of length at most $5$.

### v3

@@ 1. Introduction
In this paper, we are interested in a small variation on these questions. More precisely, we fix only the bare minimum (so one subset, or a pair of points) and we do not aim to find a trivial stabilizer, since this would be impossible in general, but to find a stabilizer that has at least some good properties. We will see that this can be done in a few interesting situations. Regarding subsets, Babai [Bab22] has proved that in every solvable permutation group there exists a set-stabilizer whose orbits have length bounded by a nonexplicit absolute constant. Answering [Bab22, Question 7a], our first main result provides the best possible constant:
@@
Let $G\leqslant\mathop{\mathrm{Sym}}(\Omega)$ be a solvable permutation group. Then there exists a subset of $\Omega$ whose setwise stabilizer has all orbits of length at most $6$. The constant $6$ is the best possible.
@@
It happens that metabelian stabilizers need not to exist in general (see Example 2.11), so Corollary 1.2 answers [Bab22, Question 7b]. Incidentally, the same example answers negatively a question in [Glu24] on the existence of supersolvable set-stabilizers. We stress that the two constants $6$ and $3$ above turned out to be a little higher than what initially expected, and better results can be obtained for primitive solvable groups. As far as we know, Theorem 1.1 is the first statement concerning stabilizers of imprimitive groups where the property involved is not preserved by a group extension, and part of the difficulty of the proof is in finding the right statement to prove by induction.
With similar ideas, we are able to show that, in every solvable permutation group, stabilizing two disjoint subsets can force the orbits to have length at most $2$. This gives a strong positive answer to [Bab22, Question 7c]. Moreover, we give some elementary proofs concerning nonsolvable groups: we provide an easier proof of a theorem of Cameron [Cam86], and answer [Bab22, Question 6a]. In particular, we prove the following:
@@
Let $G$ be any permutation group on a finite set. Then there exists a setwise stabilizer whose nonalternating composition factors have order bounded by an absolute constant.
This is new even using the Classification of Finite Simple Groups.
We come to the question of fixing two points. Extending the work of Seress [Ser96], Burness [Bur21] showed that in a primitive group with solvable stabilizer, there always exist $5$ points whose pointwise stabilizer is trivial. The hypothesis of having a solvable stabilizer is natural, if we aim to structural results on the stabilizer of two points. For primitive solvable groups, in [LS24] the authors observed that there exist two points whose pointwise stabilizer has bounded derived length. On the other hand, motivated by various applications, Li and Zhang [LZ11] have classified the solvable maximal subgroups of the almost simple groups. Combining their work with some properties of the finite simple groups, we remark the following:
Theorem 1.4.
Let $G$ be an almost simple group, and let $H<G$ be a maximal subgroup. If $H$ is solvable, then the derived length of $H$ is at most $10$. The constant $10$ is attained only by the solvable maximal subgroups of the Fischer group $Fi_{{}_{23}}$.
@@
The proof of Theorem 1.5 is an application of Corollary 1.2, and it is perhaps surprising that, to answer a question about pointwise stabilizers of primitive nonsolvable groups, one has to deal with set-stabilizers of arbitrary solvable permutation groups. The idea, usually applied in the context of imprimitive linear groups, is using a partition to construct a vector with small centralizer (see [MW04, Sec. 4]). Our argument involves a double coset decomposition and a neat bound gives the constant $13$. It is likely that a more in-depth study of almost simple groups leads to an improvement, but the best possible constant is not even known for solvable groups [LS24].
The paper is organized as follows. In Section 2 we study set-stabilizers, proving Theorem 1.1 and the other related results. In Section 3, we prove Theorems 1.4 and 1.5. The methods are a mixture of group theory and combinatorics.
Acknowledgments: I am grateful to S. Eberhard for many useful conversations. After the public announcement of a first version of this paper, I received precious comments from D. Gluck, H.Y. Huang and S.V. Skresanov. The first asked for a structural result in the spirit of Corollary 2.14 (see also [Glu24]). The second noticed that the (relatively high) constant $10$ in Theorem 1.4 is attained by $Fi_{{}_{23}}$, and that the constant in Theorem 1.5 is at least $5$ for almost simple groups. The third brought the article [Bab22] to my attention, and helped me with the computer calculations in Lemma 2.7. After this, I did a substantial rewrite of the paper and had useful conversations with L. Babai. Finally, I am thankful to L. Pyber for the many ideas he has given me in these years.
2. Breaking symmetry with set-stabilizers
@@
When $G$ acts on a finite set $\Omega$, then $G$ acts naturally on the power set $\mathcal{P}(\Omega)$. If $\Delta\subseteq\Omega$, we write $C_{G}(\Delta)$ for the pointwise stabilizer, and $Stab_{G}(\Delta)$ for the setwise stabilizer. For each $g\in G$, it is easy to see that the stabilizers of $\Delta$ and $\Delta^{g}$ are conjugate subgroups. As usual, we say that $\Delta$ lies in a regular orbit (i.e., the orbit $\Delta^{G}$) if $Stab_{G}(\Delta)=1$. Of course, as subgroups of $G$, the stabilizers still act on $\Omega$.
Definition.
For $i\geq 1$, we say that $\Delta\subseteq\Omega$ lies in a $i$-regular orbit if all the orbits of $Stab_{G}(\Delta)$ on $\Omega$ have length at most $i$.
The $1$-regular orbits are precisely the genuine regular orbits. Moreover, unless explicitly stated otherwise, the $6$-regular orbits will also be called good orbits. (This is motivated by Theorem 1.1.) We note that, by definition,
@@
We point out that $f(b)=O(\log b)$ [Dix68]. When combined with [Ser96], this gives that in every primitive solvable group there exists a set-stabilizer of derived length at most $f(4)=3$. We are going to see that a little more is true.
@@ 2.1. Primitive solvable groups
$\dfrac{1^{1}}{1_{1_{1}}}$ GAP Id Regular orbits on $\mathcal{P}(\Omega)$ $2$-regular orbits on $\mathcal{P}(\Omega)$ $3$-regular orbits on $\mathcal{P}(\Omega)$ Good orbits on $\mathcal{P}(\Omega)$
$\ell_{2,1}$ $\ell_{2,2}$ $\ell_{2,3}$ $\ell_{2,6}$
@@
As an example to help reading Table 1, we remark that PrimitiveGroup(4,2) is $\mathop{\mathrm{Sym}}(4)$. There is no regular orbit of $\mathop{\mathrm{Sym}}(4)$ on $\mathcal{P}(\{1,2,3,4\})$, the stabilizer of $\{1,2\}$ has two orbits of length $2$, and the stabilizers of $\{1\}$ and $\{1,2,3\}$ have orbits of length $3$. We also note that PrimitiveGroup(8,2) is $\mathrm{A\Gamma L}_{1}(8)$, while PrimitiveGroup(9,7) is $\mathop{\mathrm{AGL}}_{2}(3)$. The following result is obtained with similar considerations:
Lemma 2.3.
Proof.
As before, using [MW04, Lemma 4.1(ii)], and the library PrimitiveGroups in [GAP4]. We observe that $\mathrm{A\Gamma L}_{1}(8)$ and PrimitiveGroup(9,6) do not have any $2$-regular orbit on $\mathcal{P}(\Omega)$, but still have some abelian set-stabilizer. ∎
We conclude this subsection with the following result:
@@
This is [MW04, Lemma 4.1(iii)]. As before, the proof works by checking the finitely many cases with $|\Omega|\leq 9$. ∎
Lemma 2.4 is sharp, as shown by $\mathop{\mathrm{Sym}}(4)$.
2.2. Good colorings
The standard argument to deal with imprimitive groups is [Glu83, Theorem 2], which was used by Gluck to show that permutation groups of odd order have a regular orbit on the power set. This does not work with a property such as “have all orbits of length at most $C$”, where $C$ is a constant. We use a different method based on good orbits and partitions that we are going to explain. A key observation, already in the lines of the proof of [Bab22, Theorem 12.1], is that studying stabilizers of partitions can also be useful to obtain results concerning stabilizers of a single subset.
We need some more definitions. An asymmetric coloring of a permutation group $G$ on $\Omega$ is a partition of $\Omega$ with the property that only the identity fixes every part. We refer the reader to [Bab22, Sec. 2] for some history concerning breaking-symmetry results and some conflicting terminology. We refine this notion as follows:
@@ 2.2. Good colorings
Let $G\leqslant\mathop{\mathrm{Sym}}(\Omega)$, and $k\geq 2$. A $k$-coloring is a coloring of $\Omega$ with at most $k$ colors. For $i\geq 1$, an $i$-asymmetric $k$-coloring is a $k$-coloring such that the stabilizer of the coloring in $G$ has all orbits on $\Omega$ of length at most $i$.
The $1$-asymmetric colorings are precisely the genuine asymmetric colorings. Unless explicitly stated otherwise, the $6$-asymmetric colorings will also be called good colorings. The total number of $k$-colorings of $\Omega$ is $k^{|\Omega|}$, but many of these are equivalent under the action of $G$. Fixed $G\leqslant\mathop{\mathrm{Sym}}(\Omega)$, for each $k\geq 2$ we write
@@
for the numbers of $G$-orbits of asymmetric, $2$-asymmetric, $3$-asymmetric and good $k$-colorings. When necessary, we write $\ell_{k,i}(G)$ to stress the group $G$ we are referring to. It is obvious that, for every $k$ and $i\leq j$,
(2.1) $\ell_{k+1,i}\geq\ell_{k,i},\hskip 28.45274pt\mbox{ and }\hskip 28.45274pt\ell_{k,j}\geq\ell_{k,i}.$
We observe that describing a $2$-coloring is equivalent to describing a subset. Therefore, $\ell_{2,1}$, $\ell_{2,2}$, $\ell_{2,3}$ and $\ell_{2,6}$ coincide respectively with the numbers of regular, $2$-regular, $3$-regular and good orbits on the power set. Restated in terms of colorings, Lemma 2.4 says that $\ell_{4,1}\geq 1$ for every primitive solvable group (one color is needed for the complement).
A portion of the proof of Theorem 1.1 relies on showing that $\ell_{6-i,i}\geq 5$ for all primitive solvable groups and each $i=1,2,3$. The following combinatorial lemma is useful to estimate $\ell_{k,i}$ for large $k$.
@@
Fix $k\geq 3$ and $i\in\{1,2,3,6\}$. If $\ell_{2,i}\geq 1$, then $\ell_{k,i}\geq{k\choose 2}$. If $\ell_{j,i}>\ell_{j-1,i}$ for some $3\leq j\leq k$, then $\ell_{k,i}\geq{k\choose j}$.
@@
The inequality $\ell_{j,i}>\ell_{j-1,i}$ means that there exists a suitable coloring that involves precisely $j$ colors. Now suppose we have such a $j$-coloring. For each subset of cardinality $j$ of the $k$ available colors, we obtain an inequivalent suitable $j$-coloring. ∎
Corollary 2.6.
Let $A$ be a primitive solvable group. Then $\ell_{5,1}(A)\geq 5$.
Proof.
Fix $A$. We have $\ell_{4,1}(A)\geq 1$ by Lemma 2.4. This means that $\ell_{2,1}(A)\geq 1$, or the inequality $\ell_{j,1}(A)>\ell_{j-1,1}(A)$ occurs at some index $j=3,4$. Considering the worst case, we obtain
$\ell_{5,1}(A)\geq\min_{j=2,3,4}{5\choose j}=5,$
where we used Lemma 2.5 with $k=5$, $i=1$. ∎
The next result concerns $3$-colorings.
@@
Let $A$ be a primitive solvable group. If $A\neq 1$, then $\ell_{3,2}(A)\geq 5$.
@@
Let $\Omega$ be the base set, so that $|\Omega|\geq 2$. By Lemma 2.2 and (2.1), we can assume $|\Omega|\leq 9$. Since we are interested in a lower bound on the number of orbits, it is enough to deal with the solvable (over)groups
@@
It is easy to see that $\ell_{3,2}$ is at least $5$ (actually $6$) in all these cases. In particular, using [GAP4], the command
@@
2.3. Imprimitive groups
We now explain the general strategy to deal with colorings of imprimitive groups. A similar idea is contained in [Ser96, Proof of Theorem 1.2], but we will go a little further. Let $A$ and $B$ be permutation groups on the finite sets $\Omega_{1}$ and $\Omega_{2}$ respectively. Let $G\leqslant B\wr_{|\Omega_{1}|}A$ act imprimitively on $\Omega=\Omega_{1}\times\Omega_{2}$. Let $x,y\geq 2$ be integers. Let $X=(X_{i})_{i=1}^{x}$ be a $x$-coloring of $\Omega_{1}$, namely $\Omega_{1}=\sqcup_{i=1}^{x}X_{i}$. Moreover, suppose we have $x$ inequivalent $y$-colorings of $\Omega_{2}$, say $Y_{1},\ldots,Y_{x}$. In particular, for each $i=1,\ldots,x$, we have that $\Omega_{2}=\sqcup_{j=1}^{y}Y_{i,j}$. The idea is to use the $y$-colorings of $\Omega_{2}$ as colors for $\Omega_{1}$. (In fact, we intentionally used the same index $i$ to count the colors in $X$ and the colorings of $\Omega_{2}$). We will construct a new $y$-coloring of $\Omega$, say $Z=(Z_{j})_{j=1}^{y}$, as follows. For each $j=1,\ldots y$, define
$Z_{j}\>=\>\{(\omega_{1},\omega_{2})\in\Omega:\omega_{1}\in X_{i}\mbox{ and }\omega_{2}\in Y_{i,j}\mbox{ for some fixed }i\}.$
It is easy to see that $(Z_{j})_{j=1}^{y}$ is actually a partition of $\Omega$.
Lemma 2.8.
The stabilizer of the coloring $Z$ in $G$ can be embedded in the semidirect product between the direct product of the stabilizers of the colorings $Y_{i}$’s in $B$ and the stabilizer of the coloring $X$ in $A$.
Proof.
First suppose $G=B\wr_{|\Omega_{1}|}A$. Since the colorings of $\Omega_{2}$ are inequivalent, the stabilizer of $Z$ is precisely the above semidirect product. The general case follows. ∎
Lemma 2.9.
Inequivalent $x$-colorings of $\Omega_{1}$ provide inequivalent $y$-colorings of $\Omega$.
Proof.
Fix $Y_{1},\ldots,Y_{x}$ colorings of $\Omega_{2}$, and let $X,X^{\prime}$ be two colorings of $\Omega_{1}$. Let $Z$ and $Z^{\prime}$ be the colorings of $\Omega=\Omega_{1}\times\Omega_{2}$ obtained from $X$ and $X^{\prime}$ respectively. If $Z$ and $Z^{\prime}$ lies in the same orbit of $B\wr_{|\Omega_{1}|}A$, then the induced colorings of $Z$ and $Z^{\prime}$ on $\Omega_{1}$ need to be in the same orbit of $A$. But these restrictions are precisely $X$ and $X^{\prime}$, and the proof follows. ∎
We are ready to prove Theorem 1.1. We actually prove the following stronger statement, that lends itself to an inductive argument:
@@ 2.3. Imprimitive groups
We work by induction on the cardinality of $\Omega$. Suppose that $\Omega_{1},\ldots,\Omega_{t}$ are the orbits of $G$ on $\Omega,$ and that we can find subsets $\Delta_{j}\subseteq\Omega_{j}$ such that $Stab_{G}(\Delta_{j})$ has orbits of length at most $6$ on $\Omega_{j}$, for each $j=1,\ldots,t$. Then $Stab_{G}(\sqcup_{j}\>\Delta_{j})$ will surely work. It follows that we can suppose that $G$ is transitive. Primitive groups can be handled with Lemma 2.2 and Table 1. Let $\Omega=\Omega_{1}\times\Omega_{2}$, with $|\Omega_{1}|\geq 2$. Let $G\leqslant B\wr_{|\Omega_{1}|}A$, with $A$ primitive on $\Omega_{1}$, and $B$ transitive on $\Omega_{2}$. By the inductive hypothesis applied to $\Omega_{2}$, we are in one of the following cases:
@@
    Since the goal is good $2$-colorings, we will use the construction above with $y=2$. In this case, set $x=5$. Let $\Delta_{1},\ldots,\Delta_{5}\subseteq\Omega_{2}$ be such that $\Delta_{1}^{B},\ldots,\Delta_{5}^{B}$ are distinct good orbits of $B$ on $\mathcal{P}(\Omega_{2})$. For each asymmetric $5$-coloring of $\Omega_{1}$, say $X=(X_{i})_{i=1}^{5}$, we construct a good orbit $\Sigma^{G}$ for $G$ as follows:
    $\Sigma\>=\>\{(\omega_{1},\omega_{2})\in\Omega_{1}\times\Omega_{2}:\omega_{1}\in X_{i}\mbox{ and }\omega_{2}\in\Delta_{i}\mbox{ for some fixed }i\}.$
    In fact, by Lemma 2.8, the set-stabilizer of $\Sigma$ in $G$ can be embedded in the semidirect product between the direct product of the set-stabilizers of the $\Delta_{i}$’s in $B$ (which have orbits of length at most $6$), and the stabilizer of the asymmetric $5$-coloring in $A$ (which is trivial). It follows from Lemma 2.9 that there are at least $\ell_{5,1}(A)$ good orbits of $G$. The proof of this part is completed by Corollary 2.6.
@@
    Here $\ell_{2,2}(B)=3$. Using the $2$-asymmetric orbits of $B$, and the $3$-asymmetric $3$-colorings of $\Omega_{1}$, arguing as before we can construct $\ell_{3,3}(A)$ good orbits of $G$ (set $x=3$, $y=2$ in the construction above). In fact, note that orbit lengths are multiplied in the imprimitive action, so the length of each orbit will be bounded by $2\cdot 3=6$. Now $\ell_{3,3}(A)\geq\ell_{3,2}(A)\geq 5$ by (2.1) and Lemma 2.7.
@@
    We can assume $B\cong\mathop{\mathrm{Sym}}(3)$, so that $\ell_{2,3}(B)=4$. Using the $3$-asymmetric orbits of $B$, and the $2$-asymmetric $4$-colorings of $\Omega_{1}$, as before we can construct $\ell_{4,2}(A)$ good orbits of $G$ (set $x=4$, $y=2$ in the construction above). The length of each orbit will be bounded by $3\cdot 2=6$. Finally, $\ell_{4,2}(A)\geq\ell_{3,2}(A)\geq 5$ by (2.1) and Lemma 2.7.
The proof is complete. ∎
The cardinality of the subset in Theorem 1.1 grows linearly with the cardinality of $\Omega$, which is necessary as shown by $B\wr K$, when $B$ has some orbit of length exceeding $6$, and the degree of $K$ grows. The bound on the derived length follows easily:
@@
In a previous version of this paper, Corollary 1.2 was obtained directly, without first bounding the length of the orbits. Indeed, the idea of proving a stronger statement by induction and gluing colorings is flexible enough to give structural results directly (see Theorem 2.16 below, for example). The next example provides a transitive solvable group where no subset gives a metabelian stabilizer:
@@
Let $B=\mathop{\mathrm{Sym}}(4)\wr_{4}\mathop{\mathrm{Sym}}(4)$ act imprimitively on $\Omega=\{a,b,c,d\}\times\{1,2,3,4\}$. It is not hard to see that $B$ has a unique orbit on the power set whose set-stabilizer is metabelian, namely $\Delta^{B}$, where
@@
The set-stabilizer $H$ of $\Delta$ in $B$ is isomorphic to $\mathop{\mathrm{Sym}}(3)\times\mathop{\mathrm{Sym}}(3)\times(C_{2}\times C_{2})\wr_{2}C_{2}$. Now let $G=B\wr_{2}\mathop{\mathrm{Sym}}(2)$ act on $\Omega\times\{\alpha,\beta\}$. If we color both $\Omega\times\{\alpha\}$ and $\Omega\times\{\beta\}$ using $\Delta$, then the resulting set-stabilizer will be isomorphic to $H\wr_{2}\mathop{\mathrm{Sym}}(2)$, whose derived length is $3$. Otherwise, we are forced to color a subset, say of $\Omega\times\{\alpha\}$, which does not provide a metabelian set-stabilizer in $B$. So there is no hope of getting a metabelian set-stabilizer in $G$.
Since $\mathop{\mathrm{Sym}}(3)\wr_{2}\mathop{\mathrm{Sym}}(2)$ is not supersolvable, neither is $H\wr_{2}\mathop{\mathrm{Sym}}(2)$, and the same example shows that supersolvable set-stabilizers need not to exist. This is an improvement on [Glu24, Example 3.4], which deals with nilpotent set-stabilizers.
The next similar construction involves $\mathrm{A\Gamma L}_{1}(8)$, and provides a transitive solvable group where no subset gives a stabilizer with all orbits of length at most $5$. We observe that, if a finite group acts on a set (even if not transitively), then the length of each orbit has to divide the order of the group. So, for example, $\mathrm{A\Gamma L}_{1}(8)$ does not have any orbit of length $5$.